How do you work the problem (9+2i)+(1-7i) ?
Answer by sdmmadam@yahoo.com(530)

(9+2i)+(1-7i) ----(1)
=9+2i+1-7i (attaching of brackets as well as removing of brackets is a free tool granted by associativity and here it is additive associativity)
=9+(2i+1)-7i (by additive associativity)
=9+(1+2i)-7i (by additive commutativity)(notice the swapping of 1 and 2i)
=(9+1)+(2i-7i) (by additive associativity)
=10-5i
Answer: 10-5i
Note: This above stepwise journey towards an otherwise oral problem is to give you an insight into the chronological steps involved algebraically in finally grouping like terms.
If you are answering an oral test then simply one glance at the given expression will register the following facts in a flash in your mind:
9 and 1 are like terms and are both positive and add up to 10 and
the two terms in i are such that one positive and the other negative and so give the sign of the larger and subtract the smaller from the larger giving -5i
so the answer is (10-5i)